OpenStudy (anonymous):

Use a sum-to-product formula to rewrite cos5w+cos2w as a product.

OpenStudy (jdoe0001):

http://www.sosmath.com/trig/Trig5/trig5/img9.gif

OpenStudy (anonymous):

\[\cos c+\cos d=2\cos \frac{ c+d }{ 2 }\cos \frac{ c-d }{ 2 }\]

OpenStudy (anonymous):

\[so: 2\cos \frac{ 5w+2w }{ 2 }\cos \frac{ 5w-2w }{ 2 }, then simplify?\]

OpenStudy (anonymous):

@jdoe0001 @surijthayer

OpenStudy (anonymous):

\[\frac{ \cos(5w)+\cos(2w) }{ 2 }\]

OpenStudy (anonymous):

is that right?

OpenStudy (jdoe0001):

\(\bf cos(5w)+cos(2w)=2cos\left(\frac{5w+2w}{2}\right)cos\left(\frac{5w-2w}{2}\right) \\ \quad \\ 2cos\left(\frac{7w}{2}\right)cos\left(\frac{3w}{2}\right)\) the denominator "2" is inside the cosine function, is not under the cosine function